Bibliography

collection

23/03-04 16:19

[armour2000a] Phillip G. Armour. The case for a new business model. Communications of the ACM, 43(8):19–22, August 2000.

[armour2000b] Phillip G. Armour. The five orders of ignorance. Communications of the ACM, 43(10):17–20, October 2000.

[aurrecoechea1998a] C. Aurrecoechea, A. T. Campbell, and L. Hauw. A survey of QoS architectures. ACM Multimedia Systems, 6(3), 1998. Lancaster University Report MPG-95-18.

[benedikt1991a] Michael Benedikt. Cyberspace: Some proposals. In Cyberspace — First Steps, chapter 7. The MIT Press, 1991.

[bernstein1996a] Philip A. Bernstein. Middleware: A model for distributed system services. Communications of the ACM, 39(2):86–98, February 1996.

[biersack1999a] Ernst Biersack and Werner Geyer. Synchronized delivery and playout of distributed stored multimedia streams. ACM Multimedia Systems, 7(1):70–90, January 1999.

[bosch1997c] Jan Bosch, Clemens Szyperski, and Wolfgang Weck. Summary of the second international workshop on component-oriented programming. In Wolfgang Weck, Jan Bosch, and Clemens Szyperski, editors, Proceedings of the Second International Workshop on Component-Oriented Programming (WCOP’97), number 5 in TUCS General Publication, pages 1–4. Turku Centre for Computing Science, September 1997.

[bosch1997d] Jan Bosch. Adapting object-oriented components. In Wolfgang Weck, Jan Bosch, and Clemens Szyperski, editors, Proceedings of the Second International Workshop on Component-Oriented Programming (WCOP’97), number 5 in TUCS General Publication, pages 13–21. Turku Centre for Computing Science, September 1997.

[buchi1997b] Martin Büchi and Emil Sekerinski. Formal methods for component software: The refinement calculus perspective. In Wolfgang Weck, Jan Bosch, and Clemens Szyperski, editors, Proceedings of the Second International Workshop on Component-Oriented Programming (WCOP’97), number 5 in TUCS General Publication, pages 23–32. Turku Centre for Computing Science, September 1997.

[bush1945a] Vannevar Bush. As we may think. The Atlantic Monthly, July 1945.

[cooper1994a] Wes Cooper. Virtual reality and the metaphysics of self, community and nature. Technical report, University of Alberta, May 1994.

[coulson1999a] Geoff Coulson. A configurable multimedia middleware platform. IEEE Multimedia, 6(1):62–76, January–March 1999.

[davies1988a] Paul Davies. Life: its nature. In The Cosmic Blueprint, chapter 7, pages 93–106. New York, 1988.

[dibbell1993a] Julian Dibbel. A rape in cyberspace. Village Voice, December 1993.

[fayad2000a] Mohamed E. Fayad, David S. Hamu, and Davide Brugali. Enterprise frameworks characteristics, criteria, and challenges. Communications of the ACM, 43(10):39–46, October 2000.

[feyerabend1975a] Paul Feyerabend. How to defend society against science. Radical Philosophy, (2):156–167, 1975.

[fjelland1985a] Ragnar Fjelland. Galilei og det heliosentriske system. In Vitenskap som livssyn, pages 47–52. Oslo, 1985.

[fjelland1985b] Ragnar Fjelland. Det mekanistiske verdensbilde. In Vitenskap som livssyn, pages 53–68. Oslo, 1985.

[gabriel1991b] R. P. Gabriel, J. L. White, and Daniel G. Bobrow. CLOS: Integrating objectoriented and functional programming. Communications of the ACM, 34(9):28–38, 1991.

[gemmell1995a] D. James Gemmell, Harrick M. Vin, Dilip D. Kandlur, P. Venkat Rangan, and Lawrence A. Rowe. Multimedia storage servers: A tutorial. Computer, 28(5):40–49, May 1995.

[goedicke1997b] Michael Goedicke and Torsten Meyer. Design and evaluation of distributed component-oriented software systems. In Wolfgang Weck, Jan Bosch, and Clemens Szyperski, editors, Proceedings of the Second International Workshop on Component-Oriented Programming (WCOP’97), number 5 in TUCS General Publication, pages 33–41. Turku Centre for Computing Science, September 1997.

[hafid1998b] Abdelhakim Hafid and Gregor V. Bochmann. Quality-of-service adaptation in distributed multimedia applications. ACM Multimedia Systems, 6(5):299–315, September 1998.

[hanson1961a] Norwood Russel Hanson. Observation. In Patterns of Discovery. Cambridge University Press, 1961.

[hondt1997b] Koen De Hondt, Carine Lucas, and Patrick Steyaert. Reuse contracts as component interface descriptions. In Wolfgang Weck, Jan Bosch, and Clemens Szyperski, editors, Proceedings of the Second International Workshop on Component-Oriented Programming (WCOP’97), number 5 in TUCS General Publication, pages 43–49. Turku Centre for Computing Science, September 1997.

[hopkins2000a] Jon Hopkins. Component primer. Communications of the ACM, 43(10):27–30, October 2000.

[jakobsen1990a] Trond Gansmo Jakobsen. Var den kopernikanske revolusjon en revolusjon i kuhns forstand? Stensil, Universitetet i Tromsø, 1990.

[karlsen2003a] Randi Karlsen and Anna-Brith Jakobsen. Transaction service management: An approach towards a reflective transaction service. In The 2nd International Workshop on Reflective and Adaptive Middleware, Middleware 2003, Rio de Janeiro, Brazil, 2003.

[kon2000a] Fabio Kon, Manuel Roman, P. Liu, J. Mao, T. Yamane, L. C. Magalhaes, and Roy H. Campbell. Monitoring, security and dynamic configuration with the dynamicTAO reflective ORB. In Proceedings of Middleware 2000, New York, USA, April 2000.

[krieger1998a] David Krieger and Richard M. Adler. The emergence of distributed component platforms. Computer, 31(3):43–53, March 1998.

[kvilhaug1991a] Terje Kvilhaug. Fornuftskritikk som ideologi. Agora, (1):288–305, 1991.

[liversidge1986a] Anthony Liversidge. Kunnskap på samlebånd. Arken, (1):29–35, 1986. Intervju med Erwin Chargaff.

[maurer2000a] Peter M. Maurer. Components: What if they gave a revolution and nobody came? Computer, 33(6):28–35, June 2000.

[mcdowell2001a] Charlie McDowell. On to C. Addison-Wesley, 2001. A primer to accompany Java by Dissection.

[mens1997b] Kim Mens, Cristina Lopes, Bedir Tekinerdogan, and Gregor Kiczales. Aspectoriented programming workshop report. In Cristina Lopes, Kim Mens, Bedir Tekinerdogan, and Gregor Kiczales, editors, Aspect-Oriented Programming Workshop at ECOOP’97, June 1997.

[omg1998b] Object Managment Group. The portable object adaptor. In The Common Object Request Broker: Architecture and Specification, chapter 9. Object Managment Group, 1998. (revision 2.2).

[omg1998c] Object Managment Group. General Inter-ORB protocol. In The Common Object Request Broker: Architecture and Specification, chapter 13. Object Managment Group, 1998. (revision 2.2).

[omg2001b] Object Managment Group. Security service specification. Technical report, Object Managment Group, March 2001. (version 1.7).

[omg2002a] Object Managment Group. Security service specification. Technical report, Object Managment Group, March 2002. (version 1.8).

[ousterhout1998a] John K. Ousterhout. Scripting: Higher-level programming for the 21st century. Computer, 31(3):23–30, March 1998.

[penny1994a] Simon Penny. Virtual reality as the completion of the enlightment. In Carl Eugene Loeffler and Tim Anderson, editors, The Virtual Reality Casebook. Van Nostrand Reinhold Computer, 1994.

[rakotonirainy1997a] Andry Rakotonirainy and Martin Chilvers. A distributed infrastructure for mobile computing. In The Sixth International Python Conference, pages 89–95, San Jose, California, October 1997.

[renesse1996a] Robbert van Renesse, Ken Birman, and Silvano Maffeis. HORUS: A flexible group communication system. Communications of the ACM, 39(4):76–83, April 1996.

[roeggen1994a] Inge Røeggen. Forelesninger i vitenskapsteori. Kompendium, Universitetet i Tromsø, 1994.

[sorensen1986a] Ernst Sørensen. Oswald spengler og nazismen. Arken, (1):22–28, 1986.

[stone1991a] Allucquere Rosanne Stone. Will the real body please stand up?: Boundary stories about virtual cultures. In Michael Benedikt, editor, Cyberspace — First Steps, chapter 6. The MIT Press, 1991.

[tranoy1986a] Tranøy. Vitenskap som handling: sannhet, saklighet og selvrealisering. In Samfunnsmakt og livsform, chapter IV. Universitetsforlaget, 1986.

[vogt1998a] C. Vogt, L. C. Wolf, R. G. Herrtwich, and H. Wittig. HeiRAT — quality-of-service mangement for distributed multimedia systems. ACM Multimedia Systems, 6(3):152– 166, May 1998.

[wright1995a] Gary R. Wright and W. Richard Stevens. TCP/IP Illustrated: The Implementation, volume 2 of Addison-Wesley Professional Computing Series. Addison-Wesley, 1995.

[yu2001b] Weihai Yu. Reflective transaction processing – motivations and issues. In NIK 2001, November 2001.